OpenStudy (anonymous):

Will the mean of the data set increase or decrease if 14 is removed from the set? What is the new mean? 21, 23, 22, 26, 43, 29, 28, 22, 14, 27

OpenStudy (anonymous):

it will increase and the new mean is 26.777

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

as long as you understand why

OpenStudy (anonymous):

yh i thought it was that i was just checking

OpenStudy (anonymous):

Let us first find the mean including 14 \[mean = \frac{21+23+ 22+ 26+ 43+ 29+ 28+ 22+ 14+ 27}{10}= \frac{255}{10} =25.5\] Let us first find the mean excluding 14 \[mean = \frac{21+23+ 22+ 26+ 43+ 29+ 28+ 22+ 27}{10}= \frac{241}{9} =26.77\] Since the mean after removing 14 is greater than before. So it will increase. @Answermehhhhh